Project

General

Profile

NDOS Commissioning Meeting 2011/02/17

Attendance: Jon, Mark, Mark, Chad, Andrew, Mat, Alena, Gary, Gavin, Vic, Denis, Kurt, Minerba

Agenda:
  1. Channel Recovery Plans
  2. DAQ Update & Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

Channel recovery (Mat):

  • APD removal will begin tomorrow (32 known APDs), and will check ribbon cables.
  • APD installation will begin tomorrow afternoon and continue into Saturday
  • Will try to get new pixel masks into DAQ/database

DSO scans:

  • Mark Mathis has scripts ready, should have some plots for Monday

Denis: can we start using the voltage regulators on the FEBs?

  • Tested in stand-alone mode, but not system wide
  • Jon will follow up with Leon for a plan
  • Mat: should do a DSO scan first

DAQ Software:

  • We are now reading/writing to the database for DCM configuration
    • Kurt is working on back-filling run config in dB
    • Jon is working on back-filling run history data in dB
  • DCM register readback (Denis):
    • yesterday we ran a test. First 30 minutes were fine, the DAQ kept up. After 30 minutes, issued a new timing sync and that caused problems.
    • Peter suggested that we try doing timing syncs when not doing status readback, and saw similar problems
    • So it looks like we have an issue with timing sync signals.
    • Possible that during a sync, some FEBs fall out of sync.
  • 3-hour wigouts were seen last Thursday
    • traced to how the TDU client was interacting with firmware, it was auto-resyncing too frequently
    • fixed, and now things seem to be more stable
    • Ron and Vic would like to see if they can reproduce a wigout on the test-stand

DCS (Chad): temperature alarms have disappeared

  • combination of cooling fans that Rick T installed
  • working on a script to generate DCS plots and publish them on a webpage
  • Jon has copied all "old" DCS data that was in the nova_hardware and nova_dev database into the nova_prod database.

Any Other Business: